# Soft-CAM: Making black box models self-explainable for high-stakes decisions

### v1

@@ 5.4 Ridge regularization improves explanation for large disease regions
Since the CXR dataset provided larger bounding boxes localizing disease regions, unlike the point-wise lesion annotations available in the fundus and OCT datasets, we computed activation precision [9, 43], which measures the proportion of the class-guided explanation that fall within the ground-truth bounding boxes, emphasizing precision by penalizing only false positives. However, it does not account for sensitivity or penalize false negatives. To address this limitation, we extended this metric to activation sensitivity (see Appendix B.2), which penalizes false negatives to better assess the explanation completeness, especially important in clinical imaging tasks where missing relevant regions can be critical, such as in multi-focal infectious diseases like pneumonia like pneumonia [37]. We further investigated how different regularization strategies affect explanation quality. While Lasso regularization promoted sparsity by shrinking some activations to zeros, ridge regularization encouraged small (but nonzero) values, resulting in denser evidence maps. To evaluate this, we trained a ridge SoftCAM model ($\lambda_{1}=0$) and compared its performance to dense and sparse SoftCAM, as well as to the post-hoc explanation methods. The ridge penalty values were selected to balance classification performance ($\lambda_{2}=7.10^{-5}$ vs. $\lambda_{2}=2.10^{-4}$ for ResNet and VGG; see Appendix E.1).
Under comparable classification performance (Acc.=$0.95$ for Ridge ResNet*, and VGG*), we found that all SoftCAM variants (dense, sparse, and ridge) generally outperformed the evaluated posthoc methods in both activation precision and activation sensitivity (Fig. 4; Appendix E.2, E.3). Specifically, sparse SoftCAM achieved the highest activation precision, while ridge SoftCAM excelled in activation sensitivity. Dense SoftCAM consistently performed in between, underscoring the importance of balancing lasso and ridge regularization via ElasticNet to adapt to varying interpretability needs.
5.5 SoftCAM provides resource-efficient and faithful explanations for multi-class tasks
Finally, we extended our method to the multi-class setting for retinal disease diagnosis. We retrained the same training setup as for the binary tasks, adjusting the output classes in the evidence layer to $5$ for DR grading (fundus dataset) and 4 for retinal disease classification (OCT dataset). Given the small size of retinal lesions, we used Lasso regularization, selecting $\lambda_{1}$ values that balanced performance (e.g. $\lambda_{1}=9.10^{-4}$ vs. $\lambda_{1}=3.10^{-6}$ for ResNet and VGG on the OCT dataset; Appendix F.1). Both dense and sparse models achieved performance comparable to their respective black-box baselines (Tab.1), with a slight improvement in Kappa on the fundus dataset when using the ResNet backbone.
As no ground-truth lesion annotations were available for the multi-class tasks, we evaluated the faithfulness of the explanations by measuring their contribution to model predictions. For correctly classified test samples, we progressively removed top-k ($k=30$) ranked patches (based on the explanation maps; see Sec.5.3) and tracked the average drop in class confidence. In both tasks, the dense and sparse SoftCAM achieved superior performance, with sparse SoftCAM yielding the lowest area under the deletion curve, indicating the highest faithfulness (see Appendix F.2, F.3).
Notably, the sparse SoftCAM produced class-wise explanations that aligned well with class model confidence, showing minimal evidence in healthy classes (Fig.5; Appendix F.4, F.5 for VGG and more examples). In the case of DR detection, a progressive disease, it is expected that images labeled with grade $x$, where $51<x<$ 5, may still exhibit features from earlier stages, consistent with explanations. Unlike post-hoc CAM-based methods, which require backpropagation or perturbation for each class, SoftCAM generates class-specific explanations during prediction in a single forward pass, making it more resource-efficient.
Figure 5: Examples of multi-class explanations using ResNet. For a severe DR example from the Kaggle dataset, the first row shows class-specific dense SoftCAM evidence map explanations, while the second presents explanations from the sparse SoftCAM.
6 Discussion
Here, we introduced SoftCAM, a straightforward yet effective approach for transforming black-box CNN models into inherently interpretable architectures. We tested SoftCAM on a diverse range of medical imaging tasks, including color fundus photographs, retinal OCT scans, and Chest X-rays for disease diagnosis. Importantly, SoftCAM-variants maintained performance comparable to that of the original CNN models for classification and generally outperformed post-hoc explainability techniques. SoftCAM produces explicit class evidence maps that directly contribute to the model’s prediction. This integration enables single forward-pass generation of explanations aligned with the classification output, resulting in resource-efficient and self-explainable CNNs.
We evaluated our method for two widely used CNN backbones: ResNet-50 and VGG-16, assessing both classification performance and explainability. Despite some differences, both ResNet and VGG models employ large receptive fields, resulting in low-resolution feature maps. Consequently, the class-evidence layer operates on coarse feature maps, producing coarse-grained explanations. In the future, we could explore the integration of SoftCAM with other standard architectures like ViT [19].
Our work presents a major step forward in the development of powerful self-explainable models, demonstrating that interpretable-by-design architectures can preserve, and in some cases even improve upon, the classification performance of state-of-the-art models by modifying standard, well-tested CNN architectures without the need for complicated additional concepts such as prototypes [15]. Beyond performance, SoftCAM provides deeper insights into the model-decision-making process, offering a powerful tool for understanding mistakes and detecting spurious correlations, without relying on widely used post-hoc explanation methods. By leveraging ElasticNet regularization, which is task-specific, user can flexibly balance localization precision and sensitivity according to their application needs. This is especially relevant for CNN-based classifiers deployed in high-stakes decision-making contexts. We hope this contribution will pave the way toward designing more accurate and interpretable CNN models, ultimately fostering trust, adoption, and integration in critical real-world settings such as in medicine.
Acknowledgments and Disclosure of Funding
This project was supported by the Hertie Foundation, the German Science Foundation (Excellence Cluster EXC 2064 “Machine Learning—New Perspectives for Science”, project number 390727645). The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting KD.
Appendix A Implementation Details
A.1 Datasets
We evaluate our approach on three publicly available medical imaging datasets spanning three different modalities: the Kaggle Diabetic Retinopathy (DR) [20], the Retinal OCT dataset [31], and the RSNA Chest X-ray (CXR) dataset [1].
  • Kaggle DR Dataset. This dataset comprises $88,702$ high-resolution retinal fundus images labeled for DR severity on a 5-point scale from 00 (No DR) to $4$ (Proliferative DR). After applying an automated quality filtering pipeline using an ensemble of EfficientNet models [55] trained on the ISBI2020222https://isbi.deepdr.org/challenge2.html challenge dataset, we retained $45,923$ images from $28,984$ subjects. The resulting class distribution was $73\%$ (class 0), $15\%$, $8\%$, $3\%$, and $1\%$. For binary classification (early DR detection), we grouped class {0} vs. {1,2,3,4}, yielding an imbalance of $73\%$ vs. $27\%$. Additionally, lesion annotations for $65$ images were obtained from [17] for evaluating the model’s explanations at localizing DR-related lesions.
  • Retinal OCT Dataset. This dataset consists of $108,315$ B-scans categorized into four classes: Drusen, Diabetic macular edema (DME), Choroidal neovascularization (CNV), and Normal. A separate test set of $1,000$ B-scans is provided. Following [16], we excluded low-resolution scans (width $\leq$ 496). As preliminary experiments showed that using the full dataset did not significantly improve performance, we subsampled the training set (by randomly removing half of the healthy images [16]) to $34,962$ scans ($8,616$ Drusen, $26,346$ Normal) for binary classification (Drusen vs. Normal), preserving the original class imbalance ($73\%$ vs $27\%$). Additionally, we used $40$ drusen-annotated B-scans from [16] to evaluate the model’s explanations at localizing drusen lesions. For the multi-class classification task, the training was randomly reduced to $17,200$ images while maintaining the original class distribution: $45\%$ Normal, $34\%$ CNV, $10\%$ DME, and $9\%$ Drusen.
Each dataset was split into training ($75\%$), validation ($10\%$), and test ($15\%$) sets, except for the Retinal OCT dataset, which followed an $80\%\text{-}20\%$ training-validation split, due to its predefined test set ( $250$ images per class). All training splits used in our experiments are provided in CSV format and publicly available via the project’s GitHub [^3] Code and CSV files are available at https://anonymous.4open.science/r/SoftCAM-E1A3/ repository.
A.2 Baseline models
The effectiveness of our method was evaluated using two widely adopted black-box CNN architectures: ResNet-50 [25] and VGG-16 [49]. These models were chosen due to their distinct architectures, such as depth, theoretical receptive field size, and classification head design, which allow for a broad assessment of our method’s generalizability. In both models, the standard classification head was replaced with our proposed convolutional evidence map layer to enable inherent interpretability. For ResNet50, we removed the global average pooling layer and final fully connected layer, substituting them with a class evidence layer consisting of $C$ convolutional filters ($1\times 1$, stride 1), where $C$ is the number of output classes. This layer directly produces class-specific evidence maps (Sec. 3.2).
For VGG-16, which uses a series of fully connected layers (FCLs) in its classifier head, each FCL was replaced by an equivalent $1\times 1$ convolutional layer. Specifically, an FCL of size $b_{1}\times b_{2}$ was transformed into a convolutional layer of size $b_{1}\times b_{2}\times 1\ times1$, preserving the original parameter count and model capacity. These architectural adjustments maintain model complexity and capacity while introducing interpretability directly into the classification mechanism.
A.3 Data preprocessing
Fundus images were preprocessed by cropping them to a square shape using a circle-fitting method as described in [38]. All datasets were then resized to $512\times 512$ pixels, except for the retinal OCT dataset, which was resized to $496\times 496$ to better match its original lower resolution. Image intensities were normalized using the mean and standard deviation computed from the respective training sets.
During training, consistent data augmentation strategies were applied across all datasets. These included flipping, rotation, random cropping, and translation, each applied with a fixed probability. For the Kaggle dataset, which contains color fundus images, additional color augmentations were introduced to improve generalization.
A.4 Training setup
All models were obtained from Torchvision and initialized with pretrained ImageNet weights. They were subsequently fine-tuned on each dataset using a consistent training setup. Following [16, 18], we employed the cross-entropy loss function and optimized model parameters using stochastic gradient descent (SGD) with Nesterov momentum (momentum factor of $0.9$). The initial learning rate was set to $1.10\text{-}3$, and a clipped cosine annealing learning rate scheduling was applied with the minimum learning rate set to $1.10\text{-}4$. Weight decay was set to $5.10\text{-}4$. Training was conducted for 70 epochs with a mini-batch size of 16 on an NVIDIA A40 GPU using PyTorch [42].
A.5 Baseline CAM-based methods
Gradient-based methods primarily differ in how gradients are aggregated to compute importance weights, while gradient-free methods mainly vary in how the weights are computed.
ScoreCAM
[56]. A gradient-free method that eliminates the need for gradient information by assessing the importance of each activation map based on its forward-pass contribution to the target class score, and produces the final output via a weighted sum of these maps.
LayerCAM
[30]. A gradient-based method that generates class activation maps by leveraging the element-wise product of ReLU-activated gradients and feature maps at any convolutional layer, enabling fine-grained, spatially precise visual explanations without requiring global average pooling.
GradCAM
[47]. A gradient-based approach that uses the gradients of the target class flowing into the final convolutional layer to produce a coarse localization map, highlighting important regions in the image by upsampling the resulting map.
Guided backpropagation (Guided BP)
[51]. A gradient-based approach that modifies the standard backpropagation process to propagate only positive gradients through positive activations, producing fine-grained visualizations that highlight features strongly activating specific neurons in relation to the target output.
Integrated Gradient (Itgt Grad.)
[53]. A gradient-based method that attributes model predictions to input features by computing the path integral of gradients along a straight-line path from a baseline to the actual input, yielding fine-grained explanations.
Appendix B Explainability metrics
B.1 Activation consistency
The activation consistency [18] quantifies how well local explanations (e.g., individual activations within explanation maps) globally reflect the disease and healthy samples across a dataset. Specifically, it measures whether the activation patterns in the explanation maps consistently reflect the underlying disease or healthy class labels.
Following [18], we evaluated activation consistency by computing the proportion of positive activations (indicative of disease evidence) in saliency maps of disease samples, and negative activations (indicative of the absence of disease) in those of healthy samples. These proportions were calculated over the test set to assess whether the heatmaps consistently highlight pathological features in diseased cases and suppress activations in healthy ones. This metric thus captures the alignment between the semantic meaning of activations and the ground truth labels, offering a dataset-level evaluation of the coherence of local explanations with the global classification objective.
B.2 Activation precision and activation sensitivity
Let $\mathcal{X}=\{\bf{X}\}_{i=1}^{n}$ denote a set of input images, =$\mathcal{M}=\{\bf{M}\}_{i=1}^{n}$ the corresponding binary segmentation masks, and $\mathcal{S}=\{\bf{S}\}_{i=1}^{n}$ the associated explanation or saliency maps generated by any method. Activation precision measures the proportion of the saliency map’s positive mass that lies within the annotated region (the segmentation mask) [9, 43]. To compute it, saliency maps are first preprocessed by setting negative values to zero while retaining all positive values. This highlights how much of the explanation signal aligns with human-annotated ground truth, effectively quantifying the precision of an explainability method. The activation precision is defined as:
$AP(\mathcal{M},\mathcal{S})=\frac{\sum_{M,S}\sum_{i,j}M_{i,j}.S_{i,j}}{\sum_{i,j}S_{i,j}}.$ (6)
@@ B.2 Activation precision and activation sensitivity
However, activation precision does not penalize false-negative (i.e. missed relevant regions). To address this, we introduce activation sensitivity, which captures the completeness of the explanation by evaluating how much of the annotated region is covered by the saliency map. The activation precision is defined as:
$AS(\mathcal{M},\mathcal{S})=\frac{\sum_{M,S}\sum_{i,j}M_{i,j}.S_{i,j}}{\sum_{i,j}M_{i,j}}.$ (7)
Unlike activation precision, activation sensitivity penalizes low saliency values within the mask. For example, if $M_{i,j}=1$ but $10<S_{i,j}<$ 1, the low activation will contribute little to the numerator, reflecting reduced confidence in that region. This makes activation sensitivity especially relevant in clinical tasks where completeness is critical, such as identifying multi-focal infectious diseases like pneumonia [37].
B.3 Top-k localization precison
Top-k localization precision [18] measures the ability of an explanation map to correctly highlight salient regions that overlap with ground-truth annotations. Specifically, it quantifies the proportion of the top-k positively activated regions within an explanation that match with annotated areas. In our implementation, each explanation map is first upsampled to the input resolution and then split into non-overlapping patches of size $33\times 3$ 33. These patches are ranked based on their average activation, and the top-k (k=30) most salient patches are selected. The precision is then computed as the fraction of these patches that overlap with the annotated ground-truth regions.
This metric can be viewed as a generalization of the pointing game metric [60], where only the single most activated region (top-1) is considered, to multiple regions, making it more suitable for medical imaging tasks. In such contexts, disease-relevant features (e.g., retinal lesions or pathological markers) are often spatially distributed across the image, rather than confined to a single localized area.
B.4 Failthfulness
@@
Faithfulness, also referred to as sensitivity or fidelity [39], is a widely used metric to evaluate how accurately an explanation reflects the model’s true decision-making process. It assesses whether the importance scores (attributions) assigned to input features correspond to the actual impact of those features on the model’s prediction.
@@
In our implementation, we focus on correctly classified samples from the test set. For each, the corresponding explanation map is upsampled to the input resolution and split into non-overlapping patches of size $33\times 3$ 33. These patches are ranked based on their mean activation values, and the top-k (k=30) most salient patches are iteratively occluded. After each occlusion step, we recorded the relative drop in the model’s confidence score for the predicted class. This process yields a deletion curve, from which we compute the Area Under the Deletion Curve (AUDC). A lower AUDC indicates a more faithful explanation, as it reflects a greater decline in model confidence when the most important regions (as indicated by the explanation map) are removed, suggesting that those regions were indeed critical to the model’s prediction.
Appendix C Effect of Lasso regularization on model performance for the binary tasks
@@
The Lasso regularization coefficient $\lambda_{1}$ in Eq. 5 controls the sparsity of the class evidence map, encouraging the model to localize disease regions with high precision. For each task, $\lambda_{1}$ was selected based on a trade-off between accuracy and AUC on the corresponding validation set, choosing the highest values for which classification performance did not degrade significantly.
@@
Figure 6: Model selection on validation sets under varying Lasso regularization strengths. The regularization coefficient $\lambda$ influences model performance, with notable effects on some datasets but minimal impact on the OCT dataset. The red markers indicate the selected $\lambda$ values, chosen to balance sparsity and classification performance.
Appendix D Additionnal Results
D.1 SoftCAM provides inherently interpretable visual explanations
@@
Figure 7: Example explanations generated by different methods from VGG-16. The first column shows disease images with reference annotations, indicated by green markers or bounding boxes. Each row, from top to bottom, corresponds to fundus, OCT, and Chest X-ray images, respectively. The next five columns present saliency maps generated by post-hoc explanation methods, gradient-free (ScoreCAM, LayerCAM) and gradient-based (GradCAM, Guided BP, Itgd Grad). The final two columns showcase our proposed inherently interpretable dense and sparse SoftCAM explanations.
Figure 8: Additional example explanations of disease images from the ResNet model.
Figure 9: Additional example explanations of disease images from the VGG model.
D.2 Activation consistency

### v2

@@ E.1 Activation consistency
corresponding to the proportion of positive and negative activations in the attribution map. Activation consistency is then defined as
$AC_{+}=$ $\frac{1}{|\mathcal{D}_{\text{disease}}|}\sum_{i\in\mathcal{D}_{\text{disease}}}\mathbbm{1}_{+}(A_{i});$ (5)
$AC_{-}=$ $\frac{1}{|\mathcal{D}_{\text{healthy}}|}\sum_{i\in\mathcal{D}_{\text{healthy}}}\mathbbm{1}_{-}(A_{i}),$ (6)
where $AC_{+}$ measures how consistently explanations highlight pathological regions in disease samples, $AC_{-}$ measures how consistently they suppress activations in healthy samples. This provides a dataset-level assessment of whether local explanations globally support the model’s classification behavior.
E.2 Top-k localization precison
Top-k localization precision donteu2023sparse measures how well an explanation map highlights clinically relevant regions that overlap with ground-truth annotations. Specifically, it computes the proportion of the top-k positively activated regions that coincide with annotated areas obtained from clinicians.
Given an explanation map, we first upsampled it to the original input resolution and split into non-overlapping patches of size $33\times 33$. Each patches is assigned a saliency score equal to the average activation within that patch. The top-k ($k\in[1,30]$) most salient patches are then selected, and the metric computes the fraction of these patches that overlap with annotated ground-truth regions. This generalizes the “pointing game” metric zhang2018top, which only considers the single most salient region (top-1), making it better suited for medical imaging tasks where disease-relevant features (e.g., retinal lesions or other pathological markers) are often spatially distributed across the image.
Formally, let $S$ denote the upsampled explanation map and let the image be partitioned into $P$ non-overlapping patches $\{P_{1},\ldots,P_{p}\}$. The average activation of each patch is
$a_{p}=\frac{1}{|P_{p}|}\sum_{(x,y)\in P_{p}}S(x,y).$
[…]
The Top-k localization precision is then defined as
[…]
A higher score indicates that the explanation consistently highlights clinically meaningful regions identified by experts.
E.3 Activation precision and activation sensitivity
Let $\mathcal{X}=\{\bf{X}\}_{i=1}^{n}$ denote a set of input images, $\mathcal{M}=\{\bf{M}\}_{i=1}^{n}$ their corresponding binary segmentation masks, and $\mathcal{S}=\{\bf{S}\}_{i=1}^{n}$ the explanation or saliency maps produced by a given method. Activation precision measures how much of the positive evidence highlighted by an explanation map falls within the annotated ground-truth region barnett2021case.
Before computing the metric, saliency maps are preprocessed by thresholding negative values to zero, $\mathbf{S}_{i}^{+}=\max(\mathbf{S}_{i},0)$, ensuring that only positive evidence contributes to the evaluation. For each sample, activation precision (AP) is defined as the proportion of positive activation mass contained inside the annotation mask:
$\mathbf{AP}_{i}=\frac{\sum_{p}\mathbf{S}_{i}^{+}(p)\textbf{M}_{i}(p)}{\sum_{p}\mathbf{S}_{i}^{+}(p)+\epsilon},$
where $p$ indexes spatial locations and $\epsilon$ is a small constant preventing division by zero. The dataset-level activation precision is then obtained by averaging over all samples:
$\textbf{AP}=\frac{1}{n}\sum_{i=1}^{n}\textbf{AP}_{i}.$ (8)
This metric captures how precisely the explanations signal aligns with expert annotation, providing a clinically meaningful measure of the quality of an explanation method.
@@ E.3 Activation precision and activation sensitivity
However, activation precision does not penalize false negatives (i.e. missed relevant regions). To address this, we introduce activation sensitivity, which captures the completeness of the explanation by computing the fraction of annotated regions that are covered by the saliency map. We defined per-sample activation sensitivity (AS) by computing the proportion of mask pixels that receive any positive activation:
$\mathbf{AS}_{i}=\frac{\sum_{p}\mathbf{S}_{i}^{+}(p)\textbf{M}_{i}(p)}{\sum_{p}\mathbf{M}_{i}(p)+\epsilon},$
The dataset-level activation sensitivity is obtained by averaging over samples:
$\textbf{AS}=\frac{1}{n}\sum_{i=1}^{n}\textbf{AS}_{i}.$ (9)
Intuitively, $\textbf{AS}_{i}$ is the fraction of the annotated region that the explanation covers with higher values meaning fewer missed regions.
Unlike activation precision, activation sensitivity penalizes weak activations inside the annotated region. For example, if $M_{i}(p)=1$ and $0<S_{i}(p)<1$, the low activation contributes only minimally to the numerator, indicating reduced confidence in that clinically relevant area. This makes activation sensitivity particularly valuable in settings where complete lesion coverage is essential.
E.4 Failthfulness
@@
Faithfulness, also referred to as sensitivity or fidelity yeh2019fidelity, evaluates how well an explanation reflects the model’s true decision-making process. It assesses whether the importance assigned to input features corresponds to their actual influence on the model’s prediction.
@@
In our implementation, we evaluated faithfulness on correctly classified test samples. Each explanation map is upsampled to the input resolution and split into non-overlapping $33\times 33$ patches, which are ranked by mean activation values. The top-k most salient patches are iteratively occluded, and after each step we record the relative drop in the model’s confidence for the predicted class. This produces a deletion curve from which we computed the Area Under the Deletion Curve (AUDC). A lower AUDC indicates a more faithful explanation, as it reflects a sharper decrease in confidence when the regions deemed important are removed, showing that these regions effectively contributed to the model’s decision.
Appendix F Sparsity regularization selection for the binary tasks
@@
The sparsity regularization coefficient $\lambda_{1}$ in Eq. 4 controls the sparsity of the class evidence map, encouraging the model to localize disease regions with high precision. For each task, $\lambda_{1}$ was selected based on a trade-off between accuracy and AUC on the corresponding validation set, choosing the highest values for which classification performance did not degrade significantly (Fig. 9).
@@
Figure 5: Model selection on validation sets under varying sparsity regularization strengths. The regularization coefficient $\lambda$ influences model performance, with notable effects on some datasets but minimal impact on the OCT dataset. The red markers indicate the selected $\lambda$ values, chosen to balance sparsity and classification performance.
Appendix G Additionnal Results
Visualizing explanations.
The evidence map generated by SoftCAM is upsampled to the input resolution for visualization. Like most CAM-based methods, such as GradCAM, ScoreCAM, and LayerCAM that operate on the final convolutional layer, SoftCAM’s explanations are limited by the resolution of the backbone (e.g., 16×16 for VGG-16/ResNet-50 with 512×512 input) due to pooling and striding, leading to lower-resolution saliency maps. However, by introducing the class evidence and classification layer directly on top of the low-resolution features and applying regularization, SoftCAM improves spatial precision. In contrast, gradient-based methods like Integrated Gradients sundararajan2017axiomatic and Guided Backpropagation springenberg2014striving produce high-resolution saliency maps by computing pixel-level gradients, which may lead to noisy maps, especially when the region of interest spans a broader area, as commonly observed in Chest X-ray images.
Comparison with other approaches.
Unlike post-hoc attribution-based approaches, SoftCAM is inherently interpretable from the classification layer and maintains performance comparable to its black-box counterpart, without a significant trade-off, even when regularization is applied to enhance explainability. Compared to donteu2023sparse, SoftCAM extends from interpretable bag-of-local models to general black-box CNN architectures and generalizes the regularization from lasso to ElasticNet, with extensive evaluations across multiple datasets using a broad range of explanability metrics. Compared to aubreville2019transferability, our method is trained end-to-end and does not require post-hoc processing, weight sharing between branches, or an additional forward pass to generate explanations.
G.1 SoftCAM provides inherently interpretable visual explanations
@@
Figure 6: Explanations generated by different methods from VGG-16. The first column shows disease images with reference annotations, indicated by green markers or bounding boxes. Each row, from top to bottom, corresponds to fundus, OCT, and Chest X-ray images, respectively. The next five columns present saliency maps generated by post-hoc explanation methods. The final two columns showcase our proposed inherently interpretable SoftCAM explanations.
Additional qualitative examples
Figure 7: Additional qualitative explanations from the ResNet model.
Figure 8: Additional qualitative explanations from the VGG model.
G.2 Activation consistency